# references.bib

### v1

@@ 4 Proofs
In this section we will prove the Theorem 1, Corollary 2, Proposition 3, and Theorem 4.
Proof of Theorem 1.
@@
$Y_{(G,\omega)}:=\sum_{\kappa}\prod_{v\in V(G)}\left(\prod_{i\in\kappa(v)}x_{i}\right)^{\omega(v)},$
where the sum ranges over all functions $\kappa:V(G)\to 2^{n}$ such that $\kappa(v)\cap\kappa(w)=\varnothing$ whenever $vw\in E(G).$
Lemma 10.
@@
$Y_{(G,\omega)}=\prod_{k\geq 1}(1+\overline{p}_{k})^{a_{V(G)}(k)}.$
@@
Since $Y_{(G,\omega)}$ allows vertices to not receive any colors, we can think of assigning each color to a subset of the vertices independently of all other colors. Specifically, for each color, we can independently choose a subset of $V(G)$ which will receive that color, and the only restriction is that the subset be an independent set. Thus, the generating series for the ways to choose which vertices receive color $i$ is given by the independence polynomial $I_{(G,\omega)}(x_{i})$. Since different colors can be assigned independently, we can compute $Y_{(G,\omega)}$ by simply multiplying over all colors, so
@@
Then by definition of $a_{V(G)}(k)$, we have
$I_{(G,\omega)}(x_{i})=\prod_{k\geq 1}(1+x_{i}^{k})^{a_{V(G)}(k)}.$
@@
$Y_{(G,\omega)}=\prod_{i\geq 1}I_{(G,\omega)}(x_{i})=\prod_{i,k\geq 1}(1+x_{i}^{k})^{a_{V(G)}(k)}=\prod_{k\geq 1}(1+\overline{p}_{k})^{a_{V(G)}(k)},$
@@
Theorem 1 then follows from the principle of inclusion-exclusion. The difference between $\overline{X}_{(G,\omega)}$ and $Y_{(G,\omega)}$ is that $\overline{X}_{(G,\omega)}$ requires all vertices to receive a nonempty set of colors. Thus, by inclusion-exclusion, we can compute $\overline{X}_{(G,\omega)}$ by adding the monomials all possible colorings without that restriction, then subtracting the monomials where a particular vertex receives no colors, then adding back the ones where a particular pair of vertices receive no colors, and so on. It follows that
$\overline{X}_{(G,\omega)}=\sum_{W\subseteq V(G)}Y_{(G,\omega)}.$
Plugging in the formula from Lemma 10 proves Theorem 1. ∎
The formula in Corollary 2 is now fairly immediate:
Proof of Corollary 2.
@@
Thus, get a $\overline{p}_{\lambda}$ term from $\prod_{k\geq 1}(1+\overline{p}_{k})^{a_{W}(k)},$ we need to choose a $\overline{p}_{k}^{i_{k}}$ term from $(1+\overline{p}_{k})^{a_{W}(k)}$ for each $k=1,2,\dots,\ell.$ By the binomial theorem,
@@
Summing over all $W\subseteq V(G)$ completes the proof. ∎
Next, we prove the recursive formula in Proposition 3:
Proof of Proposition 3.
@@
Expanding the left side using the power series $\log(1+x)=\sum_{i\geq 1}\frac{(-1)^{i+1}x^{i}}{i},$ we get
@@
Summing over all values $d\mid k$ and then subtracting the resulting sum from the right side coefficient $[t^{k}]\log(I_{G|_{W},\omega}(t))$ gives the formula in Proposition 3. ∎
Finally, we prove that these coefficients are all integers:
Proof of Theorem 4.
@@ Acknowledgements
The author thanks Oliver Pechenik for suggesting the problem, and she thanks Oliver Pechenik, Karen Yeats, and Sophie Spirkl for providing helpful comments. She was partially supported by the Natural Sciences and Engineering Research Council of Canada (NSERC) grant RGPIN-2022-03093.

### v2

@@ 4 Proofs
In this section we will prove Theorem 1.1, Corollary 1.2, Propositions 1.3 and 1.4, and Corollary 1.5.
Proof of Theorem 1.1.
@@
$Y_{(G,\omega)}:=\sum_{\kappa}\prod_{v\in V}\left(\prod_{i\in\kappa(v)}x_{i}\right)^{\omega(v)},$
where the sum ranges over all functions $\kappa:V\to 2^{n}$ such that $\kappa(v)\cap\kappa(w)=\varnothing$ whenever $vw\in E(G).$
Lemma 4.1.
@@
$Y_{(G,\omega)}=\prod_{k\geq 1}(1+\overline{p}_{k})^{a_{V}(k)}.$
@@
Since $Y_{(G,\omega)}$ allows vertices to not receive any colors, we can think of assigning each color to a subset of the vertices independently of all other colors. Specifically, for each color, we can independently choose a subset of $V$ which will receive that color, and the only restriction is that the subset be an independent set. Thus, the generating series for the ways to choose which vertices receive color $i$ is given by the independence polynomial $I_{(G,\omega)}(x_{i})$. Since different colors can be assigned independently, we can compute $Y_{(G,\omega)}$ by simply multiplying over all colors, so
@@
Then by the definition of $a_{V}(k)$, we have
$I_{(G,\omega)}(x_{i})=\prod_{k\geq 1}(1+x_{i}^{k})^{a_{V}(k)}.$
@@
$Y_{(G,\omega)}=\prod_{i\geq 1}I_{(G,\omega)}(x_{i})=\prod_{i,k\geq 1}(1+x_{i}^{k})^{a_{V}(k)}=\prod_{k\geq 1}(1+\overline{p}_{k})^{a_{V}(k)},$
@@
Theorem 1.1 then follows from the principle of inclusion-exclusion. The difference between $\overline{X}_{(G,\omega)}$ and $Y_{(G,\omega)}$ is that $\overline{X}_{(G,\omega)}$ requires all vertices to receive a nonempty set of colors. Thus, by inclusion-exclusion, we can compute $\overline{X}_{(G,\omega)}$ by adding the monomials all possible colorings without that restriction, then subtracting the monomials where a particular vertex receives no colors, then adding back the ones where a particular pair of vertices receive no colors, and so on. It follows that
$\overline{X}_{(G,\omega)}=\sum_{W\subseteq V}Y_{(G,\omega)}.$
Plugging in the formula from Lemma 4.1 proves Theorem 1.1. ∎
The formula in Corollary 1.2 is now fairly immediate:
Proof of Corollary 1.2.
@@
Thus, to get a $\overline{p}_{\lambda}$ term from $\prod_{k\geq 1}(1+\overline{p}_{k})^{a_{W}(k)},$ we need to choose a $\overline{p}_{k}^{i_{k}}$ term from $(1+\overline{p}_{k})^{a_{W}(k)}$ for each $k=1,2,\dots,\ell.$ By the binomial theorem,
@@
Summing over all $W\subseteq V$ completes the proof. ∎
Next, we prove the recursive formula in Proposition 1.3:
Proof of Proposition 1.3.
@@
Expanding the left side using the power series $\log(1+x)=\displaystyle{\sum_{i\geq 1}\frac{(-1)^{i+1}x^{i}}{i}},$ we get
@@
Summing over all values $d\mid k$ and then subtracting the resulting sum from the right side coefficient $[t^{k}]\log(I_{G|_{W},\omega}(t))$ gives the formula in Proposition 1.3. ∎
Now we prove the alternative recurrence from Proposition 1.4:
Proof of Proposition 1.4.
Assume we know $a_{W}(1),\dots,a_{W}(k-1)$ and want to find $a_{W}(k).$ In the product
$\prod_{j\geq 1}(1+t^{j})^{a_{W}(j)}=I_{(G|_{W},\omega)}(t),$
the coefficient $[t^{k}]$ of $t^{k}$ needs to be the same on both sides. On the right side, the coefficient is $[t^{k}]I_{(G|_{W},\omega)}(t).$ On the left side, since
$(1+t^{k})^{a_{W}(k)}+1+a_{W}(k)\cdot t^{k}+\dots,$
the $(1+t^{k})^{a_{W}(k)}$ factor contributes $a_{W}(k)$ to $[t^{k}].$ No factor $(1+t^{j})^{a_{W}(j)}$ with $j>k$ can contribute to $[t^{k}].$ Each remaining contribution come from some partition $\lambda=\ell^{i_{\ell}}\dots 2^{i_{2}}1^{i_{1}}\vdash k$ and corresponds to taking a $(t^{j})^{i_{j}}$ from the $(1+t^{j})^{a_{W}(j)}$ factor for each $j=1,2,\dots,\ell.$ By the binomial theorem, the coefficient of $(t^{j})^{i_{j}}$ in $(1+t^{j})^{a_{W}(j)}$ is $\binom{a_{W}(j)}{i_{j}},$ so the overall coefficient of $(t^{\ell})^{i_{\ell}}\dots(t^{2})^{i_{2}}(t^{1})^{i_{1}}$ is
$\binom{a_{W}(\ell)}{i_{\ell}}\dots\binom{a_{W}(2)}{i_{2}}\binom{a_{W}(1)}{i_{1}}.$
We then subtract all those coefficients from $[t^{k}]I_{(G|_{W},\omega)}(t)$ to give the formula in Proposition 1.4. ∎
As noted in §1, Corollary 1.5 about the integrality of the $\overline{p}$-coefficients is an immediate consequence of Corollary 1.2 and Proposition 1.4, but we will now show that an alternative proof can also be given by considering the transition formula between the $\overline{p}$-basis and another basis defined in [kromatic2023]: the monomial symmetric function $m_{\lambda}$ is
where the sum is taken over all ordered tuples of $k$ distinct indices $n_{1},\dots,n_{k}\in\mathbb{Z}_{>0}$. The augmented monomial symmetric function $\widetilde{m}_{\lambda}$ is
$\widetilde{m}_{\lambda}:=i_{1}!\dots i_{\ell}!\cdot m_{\lambda},$
$\overline{\widetilde{m}}_{\lambda}:=\overline{X}_{K_{\lambda}},$
Theorem 4.2 (Crew, Pechenik, and Spirkl [kromatic2023, Proposition 3.4]).
The $\overline{\widetilde{m}}$-expansion for $\overline{X}_{G}$ is
We can now use this to give our second proof of Corollary 1.5:
Alternative proof of Corollary 1.5.
@@ Acknowledgements
The author thanks Oliver Pechenik for suggesting the problem, and she thanks Oliver Pechenik, Karen Yeats, Sophie Spirkl, and the anonymous reviewers for providing helpful comments. She was partially supported by the Natural Sciences and Engineering Research Council of Canada (NSERC) grant RGPIN-2022-03093.